Property worth 185,000.00 can be purchased for 5​% down and quarterly mortgage payments of $3500.00 for 25 years. What effective annual rate of interest is​ charged?
Answer #1

Loan Amount = 185,000(1 - 0.05) = $175,750

Calculating APR,

Using TVM Calculation,

I = [PV = 175,750, PMT = -3,500, N = 100, FV = 0]

I = 6.29%

EAR = (1 + 0.0629/4)4 - 1

EAR = 6.44%